OpenStudy (anonymous):

3x - y = 7 6x - y = 10 When solving this system of equations by elimination, which could be the resulting equation when a variable has been eliminated?

OpenStudy (rock_mit182):

first you have to get rid of a variable: 1) 3x-y = 7 2) 6x -y =10 multiplying by -1 the second equation, we have: 3x - y = 7 -6x + y = -10 now adding them up -3x +0 = -3

OpenStudy (rock_mit182):

@calvinbenning now solve for x, and then replace the value of x on the equation 1) or 2) to find the value of Y

OpenStudy (rock_mit182):

Do you know how to do that ?